An athlete whirls a 6.86kg hammer tied to the end of a 1.3m chain im a simple horizontal circle.Where you should ignore any vertical deviations.The hammer moves at the rate of 0.844 rev/s. What is the centripetal acceleration of the hammer?
1
Expert's answer
2018-12-12T09:57:52-0500

Answer on Question 83678, Physics, Mechanics, Relativity

Question:

An athlete whirls a 6.86 kg hammer tied to the end of a 1.3 m chain in a simple horizontal circle. Where you should ignore any vertical deviations. The hammer moves at the rate of 0.844 rev/s. What is the centripetal acceleration of the hammer?

Solution:

The centripetal acceleration of the hammer can be found from the formula:

"a_c=v^2\/r,"

here,

"v"

is the linear velocity of the hammer,

"r=1.3 m"

is the radius of the circular path.

From the other hand, the relationship between the linear velocity

"v"

and angular velocity

"\u03c9"

of the rotating object is given by the formula:

"v=\u03c9r."

Finally, substituting this formula into the first equation, we get:

"a_c=v^2\/r=(\u03c9r)^2\/r=\u03c9^2r."

Let's substitute the numbers:

"a_c=(0.844 rev\/s * 2\u03c0 rad\/1 rev)^2*1.3 m=36.5 m\/s^2."

Answer:

"a_c=36.5 m\/s^2"

.